Reprocessing and Production Team Meeting - July 14 4PM CERN, 9 FNAL time


Vidyo Link

Attending

  • FNAL: Gaston, SeangChan, Jorge, Jen, Matteo
  • CERN : Paola, Alan, JR
  • Korea :

Personnel

  • Jen Vacation July 25-29
  • Workshop Week of Aug 16-19

News - Dima

  • Workshop week, who all is going? Agenda???? Travel plans? Let's make the best use of when we are all together!

Top issues affecting production

  • stuck workflows - stuck in running closed -Alan Update! New patches in, is everything OK now?
    • there was a bug in WmStats it was fixed yesterday, Alan is going through and cleaning things out through the agents, everything should be moving again by tomorrow.
    • need to go through the workflows sitting in acquired need to be looked at - Paola will take a look
  • Vienna put into drain with merge issues - Do we have a ticket?
    • network timeouts
  • Tails old work in system how are we going to attack, upped priority of some of the old workflows
  • workflows that are rejected on our end but still in submitted on the requestor end
  • rejected/aborted workflows that were not reported in elog: https://cms-logbook.cern.ch/elog/Workflow+processing/24957
    • JR already took care of resubmitting these, but there are more
    • Until the workshop we need to use the tools we have and make sure you elog when you reject/abort workflows so we can keep track of what is going on.

Site support -

Date Site Into the Waiting Room Out of the Waiting Room Into the morgue Out of the morgue
2016-07-07 00:00:01 T2_GR_Ioannina       x
2016-07-07 00:00:01 T2_ES_IFCA x      
2016-07-13 00:00:01 T2_BE_IIHE x      
2016-07-14 00:00:01 T2_ES_IFCA   x    

  • Keep Estonia and Vienna in manual drain until we hear back from sites that problems fixed.

Transfer Team

  • transfers from T0 to T1 has improved, 5GB/sec
  • transfers from INR tape was down for the last 5 days or so but now they are better
  • transfers that were stuck from FNAL are doing better now

Workflows

Agent Issues

  • Krista wants machines for condor/osg/kernel upgrades and reboot we need to let her know when she can do her work.

Agent redeployment

  • We are still waiting for agents to finish draining so we can do the redeploy
  • 309 may go up will hopefully go up with new version
  • 308 still needs work, keep our little paws off it it Alan is working on it.

RequestMgr2 Migration

  • we need to work ACDC's still and and see if it is functional for us, We should take a look and give feedback, this is for making the ACDC's via the web, and via the makeAllACDC's script

Merging Scripts

  • a couple changes to ReqMgr2 cloning, this should be merged in

RelVal Andrew

AOB

-- JenniferAdelmanMcCarthy - 2016-07-14

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2016-07-14 - JenniferAdelmanMcCarthy
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    CMSPublic All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback